OpenStudy (accidentalaichan):

Choose the correct description of the graph of the compound inequality

OpenStudy (accidentalaichan):

A number line with an open circle on -2, shading to the left, and a closed circle on 5, shading to the right. A number line with a closed circle on -2, shading to the left, and an open circle on 5, shading to the right. A number line with an open circle on -2, a closed circle on 5, and shading in between. A number line with a closed circle on -2, an open circle on 5, and shading in between.

OpenStudy (accidentalaichan):

\[x - 2 > -4 and 3x \le 15\]

OpenStudy (anonymous):

x - 2 > -4 x > -4 + 2 x > -2 3x <= 15 x <= 15/3 x <= 5 open circle on -2, closed circle on 5, shading in between